# HumanVBench: Exploring Human-Centric Video Understanding Capabilities of MLLMs with Synthetic Benchmark Data

### v1

@@ 8.1 Emotion Perception
Emotion Recognition in Conversation is reconstructed from a subset of the MELD dataset. Its purpose is to determine the emotion category of the current speaker, requiring analysis of the conversation context and the scene depicted in the video recording. This task evaluates the model’s ability to integrate multi-modal emotion understanding when speech text is available as contextual information. An example is shown in Figure 11.
Figure 11: Example of Emotion Recognition in Conversation task.
@@ 9 Annotations Details and Examples in Human-Centric Annotation Pipeline
For the in-the-wild videos collected from Pexels, we first apply splitting and filtering operations. Specifically, we begin by utilizing the video_resolution_filter, video_aesthetics_filter, and video_nsfw_filter operators to select videos that meet the following criteria: a resolution of at least 1280 in width and 480 in height, acceptable aesthetics, and appropriate content. Next, the video_split_by_scene_mapper is used to split the videos into scenes. The resulting clips are then filtered using the video_duration_filter to exclude clips shorter than 1 second and the video_motion_score_filter to remove static videos. These steps utilize existing operators in Data-Juicer, with parameters set to their default values except for the video_motion_score_filter, where the minimum motion score is set to 1.2. After completing these foundational steps, we apply the video_face_ratio_filter with a threshold of 0.65 to retain videos containing people. These videos are then processed using a series of mappers to generate fine-grained, multi-modal, human-related annotations.
@@ 10.1 Construction Details of 6 Descriptive Human-Centric Questions
10.2 Construction Details of 11 Closed-Ended Human-Centric Questions
@@
Emotion Recognition in Conversation: This task is reconstructed from the MELD test set and is a 7-class classification task. Specifically, for the seven emotions–neutral, joy, sadness, surprise, anger, disgust, and fear– and 30 samples are randomly selected for each category as questions. The historical dialogue is limited to no more than three sentences (including the current utterance), with each utterance presented in the format: “Speaker - utterance”. The specific question format is shown in Figure 11. The selected videos correspond to the current utterance for which the emotion need to be judged, and are taken directly from the videos corresponding to each utterance in MELD.
@@
Table 4: Annotation labels used in the construction process of 17 tasks
@@
Labels ER ETA AR EIC ERC T2H H2T HC ATD BTA BCA AST TSA AVSM ASD AVAD SCM
@@ 10.3 Details of Human Efforts for HumanVBench
We employed two professional full-time AI data annotators and invited two graduated-level volunteers to participate in the construction and evaluation of HumanVBench. They collaborate to complete a series of tasks, with two main annotators participating in the full data of each task and two volunteers participating in the sampled data. Inconsistencies will be identified and resolved (for example, through discussion or majority voting to reach a consensus) to ensure high quality. Specifically, their tasks include the following four parts.
@@
Cross-Verification: After completing the above steps, we obtained 17 usable tasks for evaluation. To further ensure the high quality of the questions, we conducted cross-verification on 16 tasks except Emotion Recognition in Conversaton task to reduce the impact of personal biases and errors on the benchmark. Specifically, the tasks were cross-assigned to the two major annotators. For each task, the annotator in this step was ensured to be different from those responsible for the Manual Verification and Correction or Human Annotation. Annotators were first required to answer the multiple-choice questions. For disputed questions where answers were marked “incorrect”, the new correct answer or option will be updated for this question, and these disputed questions were reassigned to another annotator for a second review. If errors persisted, both annotators discussed and agreed on a unified answer to serve as the final ground truth.

### v2

@@ 11 Annotations Details and Examples in Human-Centric Annotation Pipeline
For the in-the-wild videos collected from Pexels, we first apply splitting and filtering operations. Specifically, we begin by utilizing the video_resolution_filter, video_aesthetics_filter, and video_nsfw_filter operators to select videos that meet the following criteria: a resolution of at least 1280 in width and 480 in height, acceptable aesthetics, and appropriate content. Next, the video_split_by_scene_mapper is used to split the videos into scenes. The resulting clips are then filtered using the video_duration_filter to exclude clips shorter than 1 second and the video_motion_score_filter to remove static videos. These steps utilize existing operators in Data-Juicer [7], with parameters set to their default values except for the video_motion_score_filter, where the minimum motion score is set to 1.2. After completing these foundational steps, we apply the video_face_ratio_filter with a threshold of 0.65 to retain videos containing people. These videos are then processed using a series of mappers to generate fine-grained, multi-modal, human-related annotations.
@@
Notably, in the Human-Centric Video Annotation Pipeline, all the models we use are based on the most advanced open-source models available. As more powerful and specialized models emerge, integrating them into our pipeline can further enhance the quality of annotations.
@@ 12.1 Construction Details of 6 Descriptive Human-Centric Questions
12.2 Construction Details of 10 Closed-Ended Human-Centric Questions
@@
Labels ER ETA AR EIC T2H H2T HC ATD BTA BCA AST TSA AVSM ASD AVAD SCM
@@
Table 7: Annotation labels used in the construction process of 16 tasks
@@ 12.3 Details of Human Efforts for HumanVBench
We employed two professional full-time AI data annotators and invited two graduated-level volunteers to participate in the construction and evaluation of HumanVBench. They collaborate to complete a series of tasks, with two main annotators participating in the full data of each task and two volunteers participating in the sampled data. Inconsistencies will be identified and resolved (for example, through discussion or majority voting to reach a consensus) to ensure high quality. The average annotation time per question is 3 minutes, totaling 10 workdays. Specifically, their tasks include the following four parts.
@@
Cross-Verification: After completing the above steps, we obtained 16 usable tasks for evaluation. To further ensure the high quality of the questions, we conducted cross-verification on 16 tasks except Emotion Recognition in Conversaton task to reduce the impact of personal biases and errors on the benchmark. Specifically, the tasks were cross-assigned to the two major annotators. For each task, the annotator in this step was ensured to be different from those responsible for the Manual Verification and Correction or Human Annotation. Annotators were first required to answer the multiple-choice questions. For disputed questions where answers were marked “incorrect”, the new correct answer or option will be updated for this question, and these disputed questions were reassigned to another annotator for a second review. If errors persisted, both annotators discussed and agreed on a unified answer to serve as the final ground truth.